# Quantum annealing versus classical solvers: Applications, challenges and limitations for optimisation problems

### v2

@@ 2.5 Hybrid computing
One of the main issues with modern quantum computing and quantum annealers is their scalability. For most industrial problems, the current number of qubits is not sufficient to run purely via quantum methods. However, hybrid workflows pave a path to tackle this issue by identifying and decomposing the problem into several sub-problems. Hereby, the strengths of both classical computers and quantum annealers for suited tasks can be utilized similarly as GPU is used for training machine learning models. D-Wave’s Leap Hybrid Solvers [39] are cloud-based solvers and are the currently leading hybrid quantum annealing algorithm. Due to its proprietary nature, it acts like a black box, however, some generally known features are summarized in the following. When identifying and decomposing the problem into several subproblems, the LeapHybridCQMSolver, designed for constrained quadratic models, automatically implements an appropriate penalty term for the constraint with suitable Lagrangian multipliers. D-Wave does not guarantee that this setting is optimal as these might differ for certain problems.
After the preprocessing, a hybrid workflow [39] is implemented utilizing classical heuristics like simulated annealing and Tabu search [55]. Simultaneously the quantum hardware is used to run smaller binary quadratic unconstrained optimisation problems with the potential to boost the classical counterparts, noting that linear binary problems can easily be transformed into QUBOs. Furthermore, the hybrid solvers also take over the task of embedding the problem onto the quantum hardware and run the QUBO partitions several times, creating a sampleset of solutions that then can be incorporated into the hybrid workflow again. As both the quantum annealing and the integrated classical methods are heuristic, the entire output is also heuristic. Thereby the solver does not guarantee the optimal solutions but returns a sampleset of several results. The entire hybrid approach is probabilistic. Finally, the exact nature of these solvers is not known, however, D-Wave enables one to create one’s own custom hybrid workflows which, if properly implemented, might even outperform the LeapHybridSolvers.
@@ 3 Case study
This section presents different case studies and their results. The exploration covers various classes of optimization problems: Integer/Binary Linear Programming (BLP) while also considering the influence of additional constraints of both linear and quadratic nature in Section 3.1 and Integer/Binary Quadratic Programming (BQP) in Section 3.4. To assess the performance of D-Wave’s Advantage system for industrial use, we benchmark the solution quality and computational time of D-Wave’s LeapHybridCQMSolver [40] against results produced by the widely adopted classical solvers CPLEX [36], Gurobi [37], and IPOPT [38] (version 3.14.12). The classical solvers are executed on a dedicated computer cluster equipped with dual 3.5 GHz Intel Xeon Gold 5115 CPUs, each featuring 10 cores and 96 GB of RAM. Noting here that for the time comparison, we measured the total run time as outputted by the cloud service, which includes the entire time it takes to return the full sample-set of on average 100 sample solutions. It should also be noted that from the sample-set, the best solution still needs to be selected. We did not consider this in the computational time to reach sufficient solutions.
@@ 3.1 Integer/Binary Linear Programming (BLP)
Initially, our focus is on exploring optimization problems characterized by variables that are exclusively either integer or binary values, with linear objective functions and constraints. We consider how solution quality and execution times vary with the number of assigned variables for the hybrid CQM solver and compare its performance with the introduced classical ones.
The problem is formulated as
@@
where $\mu_{i}$ are random constants between (0,1) and $x_{i}\in\{0,1\}$ are binary variables. Constraint (Equation 10) restricts the sum over all binary variables to equal a fixed constant $C$. We first run the solvers with 10,000 to 100,000 binary variables. A comparison of the solutions and the computational time for this range of binary variables is shown in Figure 1(a). It can be seen that D-Wave can find the same solution as both CPLEX and Gurobi, while IPOPT starts diverging for an increasing number of variables and outputs negative objective function values. IPOPT is designed to handle large-scale nonlinear problems [38] and is not very suited for integer variables. As IPOPT approximates the binary variables with continuous ones with very small bounds around zero and one, the derived objective function value with these numbers can score a lower value than what is actually true. We excluded IPOPT’s objective function values in Figure 1(b) however, we kept this solver throughout this work as it showed important results in Section 3.4.
@@
Looking at Figure 1(a), D-Wave stays relatively constant in computation time of about 6s up to 38,000 variables, from whereon it increases more drastically when the number of variables is increased further. Viewing the objective function value, we see that IPOPT behaves similarly to the other solvers for a low number of variables, but starts diverging from the rest as the number of variables increases. It can also be noted that the computational time of IPOPT increases rapidly compared to the other classical solvers. CPLEX and Gurobi start below 1s and gradually increase with an increasing number of variables.
D-Wave finds the same optimal solution as Gurobi and CPLEX, resulting in three perfectly overlapping lines. The observed sawtooth pattern of the solutions is due to the randomness of the $\mu$values as they are newly drawn for each run with a specific seed. When comparing computational time in Figure 1(b), the hybrid solver requires significantly more time as it approaches its scalability limit. This behaviour could be explained by the need for an extremely high amount of partitioning into subproblems as the number of variables increases [34]. Although the optimal solution has been found, D-Wave’s hybrid solvers show no computational advantage here yet.
@@ 3.2 Increasing the complexity by adding more constraints
Next, the assessment focused on determining whether D-Wave could continue to find the optimal solution at its scalability limit when increasing the complexity of the BLP (Equation 9) by adding additional constraints. As the constraints are added as penalty terms, we want to see if the solution quality can be maintained even though the computational time is significantly higher. We created some arbitrary although feasible constraints listed below.
@@
The objective function value and computational time averaged over 5 runs together with the minimum and maximum deviation as vertical lines are shown in Figure 2(b).
@@
We observe that the solution quality stagnates with more constraints and in particular for the 5th added constraint a significant variability in the result with an approximate increase up to a factor of five. D-Wave’s computational time is significantly higher reaching above 600s on the fully constrained problem. Furthermore, we ran the same problem with just 10,000 variables where D-Wave was able to find the optimal value with all 5 constraints added in every one of the 5 runs. This is a clear indication that D-Wave’s CQM hybrid solver can not ensure optimal solutions at its variable limit. Although no thorough study on the best setting for the Lagrange multipliers was conducted, minor adjustments to their values showed no improvements, supporting our conclusion that D-Wave fails to provide optimal solutions at its variable limit with an increasing amount of constraints.
@@ 3.3 Integer/Binary Linear Programming with quadratic constraints
Figure 3: Mean objective function value (top) and computational time (bottom) comparison between two classical solvers and D-Wave’s hybrid solver for the BLP with a single quadratic constraint. In (a) the complexity scales with the number of variables. In (b) the variables is fixed to 500 whereas the complexity scales with increasing $C$ as stated in Equation 10. All results are averaged over 5 runs, and the vertical lines show the minimum and maximum deviation from the mean value. Gurobi’s solve time is significantly larger than all other solvers and is thereby excluded here for visibility.
In the first scenario, we observe that D-Wave finds worse solutions than both CPLEX and IPOPT whereas IPOPT’s value ranks lower due to the approximate representation of the binary variables as already mentioned in the previous section. The solution obtained by CPLEX corresponds to the actual optimal solution from which D-Wave is continuously about a factor of two off. Nonetheless, D-Wave does solve the problem faster than IPOPT from 500 variables and then CPLEX at 600 variables onwards. Although D-Wave didn’t find the optimal solution a single time, the computational time is unaffected here by the increase of variables.
We experience similar behaviour when stepwise increasing the $C$-value of the constraint. D-Wave is again slightly off the optimal solution however, it stays relatively constant in computational time and solves almost consistently faster than IPOPT. Although the solution is suboptimal, it can be seen in this figure that also for quadratic constraints with a linear objective function, the hybrid solver is not able to find the optimal solution in any run. We note that we did not adjust the Lagrange multipliers and kept the ones chosen by default by the hybrid solver. We do not rule out that a customised setting of the parameters could be beneficial towards the solution quality.
@@ 3.4 Integer/Binary Quadratic Programming (BQP)
To benchmark the different solvers for this BQP, we increase the complexity by incrementally increasing the value of the constant $C$. Thereby, the complexity of the task is increased as the feasible solution space increases. This is preferred over a further increase of $N$ to avoid an unnecessarily large memory usage as the number of terms in the objective function increases with $N^{2}$.
For the comparisons in this section, we choose a sufficiently large problem with $N=500$ although different sizes were tested, showing the same tendencies. The results are shown in Figure 4.
Figure 4: Comparison of the objective function value (a) and computational time (b) between D-Wave and the previously introduced classical solvers for 500 binary variables. The results are averaged over 5 runs. The run time is limited to 1000s, resulting in CPLEX not finding the optimal solution for increasing $C$.
Considering both solution and run time we observe a clear computational advantage of the hybrid solver to CPLEX. We cut CPLEX off after 1000s of runtime, resulting in sub-optimal solutions when the solution space increases. Gurobi performed even worse than CPLEX for this problem and was left out for visibility reasons. While IPOPT shows comparable solutions and run time to D-Wave, the solutions are still not optimal. D-Wave finds consistently better solutions in less computational time. To more clearly visualise this, we show the relative difference in solution quality between D-Wave’s hybrid solver and the objective function value of both IPOPT and CPLEX in Figure 5.
Figure 5: Mean relative difference of D-Wave’s hybrid solution and those supplied by CPLEX and IPOPT. The vertical lines indicate the difference between CPLEX’s and IPOPT’s mean solution to the maximum and minimum solution provided by D-Wave within all 5 runs.
We see from Figure 5 that CPLEX performs well in the lower complexity region and that D-Wave does not find exclusively better solutions than CPLEX here, however, due to the capped computational time D-Wave outperforms consistently for higher values of $C$. While IPOPT’s relative difference becomes increasingly small, CPLEX, with a capped solve-time of 1000s, seems to settle at a deviation of $\sim 25\%$ from the optimal solution. This shows that D-Wave’s hybrid solvers are a suitable alternative for BQPs, even showing a computational advantage for this specific case. However, in the field of optimisation, there is only a very limited amount of such problems that correspond to these pure BQPs.
@@ 4 The unit commitment problem – An energy system application
With the release of the LeapHybridCQM solver [40], D-Wave opened itself up to a wider range of problems, such as Mixed Integer Linear Problems (MILP), with a wide spread of applications in many industrial sectors. We chose to benchmark D-Wave’s current performance in a real-world Unit Commitment problem which encompasses a large set of relevant optimisation problems in the energy sector and has high applicability to general power systems. It focuses on minimizing the overall cost of power generation within a specified time frame by optimizing the scheduling of generating units. As an MILP, it contains both continuous and binary variables, whilst remaining purely linear in the objective function and constraints. The problem is tested at three different scales as defined in Table 1. The objective function of a unit commitment model is as follows:
min $\text{min }\sum_{g\in G}\sum_{t\in T}\left(c_{g}(t)+CP_{g}^{l}\,u_{g}(t)+\sum_{s=1}^{S_{g}}\left(CS^{s}_{g}\delta^{s}_{g}(t)\right)\right),$ (15)
where $c_{g}(t)$ is the cost of the power produced by thermal generator $g$ above the minimal requirement at time $t$, $CP_{g}^{l}$ represents the cost of operating at generation point $l$ for generator $g$ and $u_{g}(t)$ is the commitment status of generator $g$ at time $t$. The final term contains information about the startup cost in category $s$ for generator $g$, denoted $CS_{g}^{s}$ and the startup in category $s$ for thermal generator $g$ at time $t$, denoted $\delta^{s}_{g}(t)$, summed over all startup categories $s\in S_{g}$, for thermal generator $g$. These three terms are summed over all periods in time $t\in T$ and over all the involved generators $g\in G$. We adjusted the model to our needs from [58] and a complete description of the problem and its constraints can be found therein.
For the full-scale problem, the default pre-set minimum run time is set by D-Wave at 20s. In this limit, the hybrid solver was not able to find a single feasible solution, shown in Figure 6(a).
We therefore solved the problem 6 times with an increased minimum solve time of 150s and and 2 times of 400s, shown in Figure 6(b).
For comparison, the near-optimal solution with a gap of 0.05% was found by Gurobi in 180s with a value of $1.23\times 10^{6}$. We can observe a clear increase in solution quality by the hybrid solver as in both cases feasible solutions were found. The best feasible solution for the 150s case yielded within 6 runs an objective function value of $5.9\times 10^{6}$, while two runs with 400s run time yielded a value of $5.7\times 10^{6}$. This is a clear indication that the minimum run time chosen by the solver itself is not optimal and has to be adjusted specifically for the problem at hand as a balance between computational time and solution quality.
@@
Next, an investigation was conducted to observe how D-Wave would perform if the problem was downsized, as stated in Table 1. First, the Reduced-1 problem with 12 time periods and 4 segments was examined, making it approximately one-fourth of the size of the initial problem. We kept the default minimum run-time setting of 10s, as this is already significantly longer than Gurobi’s solve time of 1.34s for the optimal solution without a gap. The solution distribution of 70 runs is shown in Figure 7(a), while the deviation of each run to the classical optimal solution is presented in Figure 7(b).
@@
Figure 6: Solution distribution from D-Wave’s LeapHybridCQM solver for the full-scale unit commitment problem. (a) shows the results of 50 runs with the solver’s default chosen run time of 20s. (b) shows the solution for the same problem with 6 runs for an increased run time of 150s and 2 runs for one of 400s. Each run supplies a sampleset with on average 100 solutions.
Figure 7: Solution distribution of the Reduced-1 unit commitment problem. (a) shows the objective function values of 70 (blue) runs given by D-Wave’s LeapHybridCQM solver with an average of 100 solutions per iteration. The feasible solutions are highlighted in orange. (b) shows the absolute difference between the optimal classical and the best feasible hybrid solution for each run for the Reduced-1 problem. Only one of the 70 runs did not obtain a feasible solution.
Although D-Wave found at least one feasible solution in every run except for one, its best solution differs by approximately a factor of 10 from the optimal.
The Reduced-2 problem is reduced even further to just 2 time periods and a single segment of linearisation. The solution distribution over 70 runs is shown in Figure 8(a) and the deviation from the optimal classical solution in Figure 8(b).
Figure 8: Solution distribution of the Reduced-2 unit commitment problem. (a) shows the objective function values of 70 (blue) runs given by D-Wave’s LeapHybridCQM solver with an average of 100 solutions per iteration. The feasible solutions are highlighted in orange. (b) shows the absolute difference between the optimal classical and the best feasible hybrid solution for each run for the Reduced-2 problem. A feasible solution was found in every iteration.

### v3

@@ 2.5 Hybrid computing
One of the main issues with modern quantum computing and quantum annealers is their scalability. For most industrial problems, the current number of qubits is not sufficient to run purely via quantum methods. However, hybrid workflows pave the way to tackle this issue by identifying and decomposing the problem into several sub-problems. Hereby, the strengths of both classical computers and quantum annealers for suited tasks can be utilised similarly to GPUs used for training machine learning models. D-Wave’s Leap Hybrid Solvers [44] are cloud-based solvers and are the currently leading hybrid quantum annealing algorithm. Due to its proprietary nature, it acts like a black box, however, some generally known features are summarised in the following. When identifying and decomposing the problem into several sub-problems, the LeapHybridCQMSolver, designed for constrained quadratic models, automatically implements an appropriate penalty term for the constraint with suitable Lagrangian multipliers. D-Wave does not guarantee that this setting is optimal as these might differ for certain problems.
After the preprocessing, a hybrid workflow [44] is implemented utilising classical heuristics like simulated annealing and Tabu search [37]. Simultaneously, the quantum hardware is used to run smaller binary quadratic unconstrained optimisation problems with the potential to boost the classical counterparts, noting that linear binary problems can easily be transformed into QUBOs. Furthermore, the hybrid solvers also take over the task of embedding the problem onto the quantum hardware and run the QUBO partitions several times, creating a sampleset of solutions that then can be incorporated into the hybrid workflow again. As both the quantum annealing and the integrated classical methods are heuristic, the entire output is also heuristic. Thereby the solver does not guarantee the optimal solutions but returns a sampleset of several results. The entire hybrid approach is probabilistic. Finally, the exact nature of these solvers is not known, however, D-Wave enables the creation of custom hybrid workflows which, if properly implemented, might outperform the LeapHybridSolvers.
@@ 3 Case study
This section presents different case studies and their results. The exploration covers various classes of optimisation problems: Integer/Binary Linear Programming (BLP), while also considering the influence of additional constraints of both linear and quadratic nature in Section 3.1, and Integer/Binary Quadratic Programming (BQP) in Section 3.4. To assess the performance of D-Wave’s Advantage system for industrial use, we benchmark the solution quality and computational time of D-Wave’s LeapHybridCQMSolver [45] against results produced by the widely adopted classical solvers CPLEX [41], Gurobi [42], and IPOPT [43] (version 3.14.12). The classical solvers are executed on a dedicated computer cluster equipped with dual 3.5 GHz Intel Xeon Gold 5115 CPUs, each featuring 10 cores and 96 GB of RAM. Noting here that for the time comparison, we measured the total run time as outputted by the cloud service, which includes the time it takes to return the full sample-set of on average 100 sample solutions. From these solutions, we first filter out the feasible set before selecting the solution with the lowest energy value. The time required for these two steps was not included in the computational time measurement for obtaining sufficient solutions. While these steps typically take less than 1s on average, their impact may only be considerable for very short total run times.
@@ 3.1 Integer/Binary Linear Programming (BLP)
Initially, our focus is on exploring optimisation problems characterised by variables that are exclusively either integer or binary values, with linear objective functions and constraints. We consider how solution quality and execution times vary with the number of assigned variables for the hybrid CQM solver and compare its performance with the introduced classical ones. The problem is formulated as:
@@
where $\mu_{i}$ are random constants between (0,1) and $x_{i}\in\{0,1\}$ are binary variables. This restricts the sum over all binary variables to equal a fixed constant $C$. We first run the solvers with 10,000 to 100,000 binary variables. A comparison of the solutions and the computational time for this range of binary variables is shown in Figure 1(a). Here, we average over 5 runs, each providing a sampleset of approximately 100 solutions from which the best feasible one is selected. As we experienced little deviation in the results we deemed higher statistics not to be necessary. Additionally, as access to quantum computers is still very limited, large statistics are not realisable. IPOPT is designed to handle large-scale nonlinear problems [43] and is not suited for integer variables. As IPOPT approximates the binary variables with continuous ones with very small bounds around zero and one, the derived objective function value with these numbers can score a lower value than what is actually true. We excluded IPOPT’s objective function values in Figure 1(b), however, we kept this solver throughout this work as it shows important results in Section 3.4.
@@
Looking at Figure 1(a), D-Wave stays relatively constant in computation time of about 6s up to 38,000 variables, from whereon it increases more drastically when the number of variables is increased further. Viewing the objective function value, we see that IPOPT behaves similarly to the other solvers for a low number of variables, but starts diverging from the rest as the number of variables increases. It can also be noted that the computational time of IPOPT increases rapidly compared to the other classical solvers. CPLEX and Gurobi start below 1s and gradually increase with an increasing number of variables. D-Wave finds the same optimal solution as Gurobi and CPLEX, resulting in three overlapping lines. The observed sawtooth pattern of the solutions is due to the randomness of the $\mu$ –values as they are newly drawn for each run with a specific seed. When comparing computational time in Figure 1(b), the hybrid solver requires significantly more time as it approaches its scalability limit. This behaviour could be explained by the need for an extremely high amount of partitioning into subproblems as the number of variables increases [35]. Although the optimal solution has been found, D-Wave’s hybrid solvers show no computational advantage.
@@ 3.2 Increasing the complexity by adding more constraints
Next, we focus on determining whether D-Wave could continue to find the optimal solution at its scalability limit when increasing the complexity of the BLP (Equation 9) by adding additional constraints. As the constraints are added as penalty terms, we want to see if the solution quality can be maintained even though the computational time is significantly higher. We consider some arbitrary, although feasible, constraints listed below.
@@
The objective function value and computational time averaged over 5 runs, together with the minimum and maximum deviation as vertical lines, are shown in Figure 2.
@@
We observe that the solution quality stagnates with more constraints and in particular for the 5th added constraint a significant variability in the result with an approximate increase up to a factor of five. D-Wave’s computational time is significantly higher reaching above 600s on the fully constrained problem. Furthermore, we ran the same problem with just 10,000 variables where D-Wave was able to find the optimal value with all 5 constraints added in every one of the 5 runs. This is a clear indication that D-Wave’s CQM hybrid solver can not ensure optimal solutions at its variable limit. Although no thorough study on the best setting for the Lagrange multipliers was conducted, minor adjustments to their values showed no improvements, supporting our conclusion that D-Wave fails to provide optimal solutions at its variable limit with an increasing number of constraints.
@@ 3.3 Integer/Binary Linear Programming with quadratic constraints
Figure 3: Mean objective function value (top) and computational time (bottom) comparison between two classical solvers and D-Wave’s hybrid solver for the BLP with a single quadratic constraint. In (a) the complexity scales with the number of variables. In (b), the variables are fixed to 500, whereas the complexity scales with increasing $C$ as stated in Equation 10. All results are averaged over 5 runs, and the vertical lines show the minimum and maximum deviation from the mean value. Gurobi’s solve time is significantly larger than all other solvers and is thereby excluded here for visibility.
In the first scenario, we observe that D-Wave finds worse solutions than both CPLEX and IPOPT whereas IPOPT’s value ranks lower due to the approximate representation of the binary variable,s as already mentioned in the previous section.
The solution obtained by CPLEX corresponds to the actual optimal solution, from which D-Wave is continuously approximately a factor of two off. Notably, although D-Wave does not find the optimal solution, the best solution found in every run is the same, resulting in no observable error bars. This indicates that D-Wave gets stuck in a local minima. Nonetheless, D-Wave solves the problem faster than IPOPT from 500 variables and CPLEX from 600 variables onwards. Although D-Wave did not find the optimal solution a single time, the computational time is unaffected by the increase of variables.
We experience similar behaviour when stepwise increasing the value of the constraint. D-Wave is again slightly off the optimal solution, however, it stays relatively constant in computational time and solves the problem almost consistently faster than IPOPT. Although the solution is suboptimal, it can be seen in Figure 3(b) that also for quadratic constraints with a linear objective function, the hybrid solver is not able to find the optimal solution in any run. We note that we did not adjust the Lagrange multipliers and kept the ones chosen by default by the hybrid solver. We do not rule out that customised settings for the parameters could be beneficial to the quality of the solution.
@@ 3.4 Integer/Binary Quadratic Programming (BQP)
To benchmark the different solvers for this BQP, we increase the complexity by incrementally increasing the value of the constant $C$. Thereby, the complexity of the task is increased as the feasible solution space increases. This is preferred over a further increase of $N$ to avoid an unnecessarily large memory usage as the number of terms in the objective function increases with $N^{2}$.For the comparisons in this section, we choose a sufficiently large problem with $N=500$, although different sizes were tested, showing the same tendencies. The results are shown in Figure 4.
Figure 4: Comparison of the objective function value (a) and computational time (b) between D-Wave and the previously introduced classical solvers for BQP with 500 binary variables. The results are averaged over 5 runs. The run time is limited to 1000s, resulting in CPLEX not finding the optimal solution for increasing $C$.
Considering both solution and run time, we observe a clear computational advantage of the hybrid solver to CPLEX. We cut CPLEX off after 1000s of runtime, resulting in sub-optimal solutions when the solution space increases. Gurobi performed even worse than CPLEX for this problem and was excluded from the figure for visibility reasons. While IPOPT shows comparable solutions and run time to D-Wave, the solutions are still not optimal. D-Wave finds consistently optimal solutions in less computational time, which again results in no error bars. To more clearly visualise this, we show the relative difference in solution quality between D-Wave’s hybrid solver and the objective function value of both IPOPT and CPLEX in Figure 5.
Figure 5: Mean relative difference of D-Wave’s hybrid solution and those supplied by CPLEX and IPOPT for theBQP. The vertical lines indicate the difference between CPLEX’s and IPOPT’s mean solution to the maximum and minimum solution provided by D-Wave within all 5 runs.
We see from Figure 5 that CPLEX performs well in the lower complexity region and that D-Wave does not find exclusively better solutions than CPLEX here; however, due to the capped computational time, D-Wave outperforms consistently for higher values of $C$. While IPOPT’s relative difference becomes increasingly small, CPLEX, with a capped solve-time of 1000s, seems to settle at a deviation of $\sim 25\%$ from the optimal solution. This shows that D-Wave’s hybrid solvers are a suitable alternative for BQPs, even showing a computational advantage for this specific case. However, in the field of optimisation, there is a limited amount of problems that correspond to these pure BQPs.
@@ 4 The unit commitment problem – An energy system application
With the release of the LeapHybridCQM solver [45], D-Wave opened itself up to a wider range of problems, such as Mixed Integer Linear Problems (MILP), with a wide spread of applications in various fields. To evaluate its performance, we benchmark D-Wave’s current performance on a real-world Unit Commitment problem, a complex optimisation challenge in the power sector that plays a crucial role in power system operation and planning. The Unit Commitment Problem involves determining the optimal scheduling of generating units—deciding when each unit should be turned on or off—over a given time horizon. The goal is to meet the forecasted load at minimum total production cost while satisfying a range of constraints, including power balance, reserve requirements, transmission limits, and individual unit constraints such as generation limits, minimum up/down times, and ramp rates [63]. For this purpose, a well-established MILP formulation of the Unit Commitment problem is chosen [64], which consists of both continuous and binary variables while maintaining a fully linear structure in its objective function and constraints. The problem is tested at three different scales as defined in Table 1. The objective function of a unit commitment model is as follows:
min $\text{min }\sum_{g\in G}\sum_{t\in T}\Bigg{(}\underbrace{c_{g}(t)+CP_{g}^{1}\,u_{g}(t)}_{i}+\underbrace{\sum_{s=1}^{S_{g}}\left(CS^{s}_{g}\delta^{s}_{g}(t)\right)}_{ii}\,\Bigg{)},$ (15)
where the objective function consists of two components: (i) the production cost and (ii) the startup cost. $c_{g}(t)$ represents the cost of the power produced by thermal generator $g$ above its minimum output at time $t$, $CP_{g}^{1}$ represents the cost of operating at the minimum power output for generator $g$, while $u_{g}(t)$ is a binary variable that indicates the commitment status of generator $g$ at time $t$. In the final term, $CS_{g}^{s}$ represents the startup cost in category $s$ for generator $g$, where $s\in S_{g}$ represents the startup categories for thermal generator $g$, with $s=1$ corresponding to the hottest state (shortest downtime) and $s=S_{g}$ to the coldest state (longest downtime). The variable $\delta^{s}_{g}(t)$ indicates the startup in category $s$ for thermal generator $g$ at time $t$. These three terms are summed over all periods in time $t\in T$ and all generators $g\in G$. For a detailed explanation of this formulation, see [65, 64].
@@
For the full-scale problem, the default pre-set minimum run time is set by D-Wave at 20s. In this limit, the hybrid solver was not able to find a single feasible solution, shown in Figure 6(a). We therefore solved the problem 6 times with an increased minimum solve time of 150s and 2 times of 400s, shown in Figure 6(b). For comparison, the near-optimal solution with a gap of 0.05% was found by Gurobi in 180s with a value of $1.23\times 10^{6}$. We can observe a clear increase in solution quality by the hybrid solver as in both cases, feasible solutions were found. The best feasible solution for the 150s case yielded, within 6 run,s an objective function value of $5.9\times 10^{6}$, while two runs with 400s run time yielded a value of $5.7\times 10^{6}$. This is a clear indication that the minimum run time chosen by the solver itself is not optimal and has to be adjusted specifically for the problem at hand, as a balance between computational time and solution quality. Next, an investigation was conducted to observe how D-Wave would perform if the problem was downsized, as stated in Table 1. First, the Reduced-1 problem with 12 time periods and 4 segments was examined, making it approximately one-fourth of the size of the initial problem. We kept the default minimum run-time setting of 10s, as this is already significantly longer than Gurobi’s solve time of 1.34s for the optimal solution without a gap. The solution distribution of 70 runs is shown in Figure 7(a), while the deviation of each run to the classical optimal solution is presented in Figure 7(b).
@@
Figure 6: Solution distribution from D-Wave’s LeapHybridCQM solver for the full-scale unit commitment problem. (a) shows the results of 50 runs with the solver’s default chosen run time of 20s. (b) shows the solution for the same problem with 6 runs for an increased run time of 150s and 2 runs for one of 400s. Each run supplies a sampleset with on average 100 solutions each.
Figure 7: Solution distribution of the Reduced-1 unit commitment problem. (a) shows the objective function values of 70 runs (blue) given by D-Wave’s LeapHybridCQM solver with an average of 100 solutions each. The feasible solutions are highlighted in orange. (b) shows the absolute difference between the optimal classical and the best feasible hybrid solution for each run for the Reduced-1 problem. Only one of the 70 runs did not obtain a feasible solution.
Although D-Wave found at least one feasible solution in every run except for one, its best solution differs by approximately a factor of 10 from the optimal. The Reduced-2 problem is reduced even further to just 2 time periods and a single segment of linearisation. The solution distribution over 70 runs is shown in Figure 8(a) and the deviation from the optimal classical solution in Figure 8(b).
Figure 8: Solution distribution of the Reduced-2 unit commitment problem. (a) shows the objective function values of 70 runs (blue) given by D-Wave’s LeapHybridCQM solver with an average of 100 solutions each. The feasible solutions are highlighted in orange. (b) shows the absolute difference between the optimal classical and the best feasible hybrid solution for each run for the Reduced-2 problem. A feasible solution was found in every run.